# A Practitioner’s Guide to Multi-turn Agentic Reinforcement Learning

### v1

@@ 6.1 How does prior model policy influence multi-turn RL training?
We distinguish between model priors (initial policies from SFT) and continual training (subsequent multi-turn online RL). This two-stage approach mirrors real-world deployment where agents learn from demonstrations before online learning. Training SFT priors on golden solutions from w2-o3-q4 TextWorld reveals that multi-turn RL with good imitation priors achieves comparable performance with dramatically fewer RL episodes. Table 7 shows that an SFT prior trained on 60 demonstrations plus 400 RL episodes achieves 85% success on w2-o3-q4, nearly matching the 88% performance of pure RL training with 5000 episodes. This represents a significant reduction in RL training data while maintaining competitive performance.
@@
Finally, we investigate cross-domain priors by training SFT models on ALFWorld demonstrations (3553 samples) then applying PPO to TextWorld, and vice versa (3000 TextWorld demonstrations to ALFWorld). We find that cross-domain priors cause rapid policy collapse during multi-turn RL, likely because demonstration biases conflict with the target environment’s action-outcome relationships, destabilizing the policy.
@@ 6.2 How do RL algorithms impact multi-turn RL training?
Understanding whether performance gains stem from our multi-turn formulation or specific algorithmic choices is crucial for establishing the generalizability of our approach. We compare PPO (heuristic policy gradient method with value function bootstrapping) against RLOO (unbiased policy gradient estimator) to isolate the contributions of our token-level credit assignment from algorithmic design decisions (Oertell et al., 2025).
Both PPO and RLOO achieve substantial improvements over base models, demonstrating that performance gains stem from our multi-turn formulation rather than PPO-specific heuristics. Table 8 shows that PPO achieves 88% success on w2-o3-q4 versus RLOO’s 51%. This gap widens on w4-o6-q8, where PPO reaches 59% success while RLOO fails completely (0%) for the 1.5B model. PPO consistently outperforms RLOO in multi-turn settings, with performance gaps increasing for complex environments. Model size also affects the gap: at 7B parameters, both algorithms perform similarly on simple tasks (97% vs 98%), but PPO maintains advantages on complex tasks (72% vs 47%). These results confirm that the performance gains are not due to PPO heuristics, evidenced by RLOO’s consistent improvements across tasks.
Table 8: Comparison of PPO and RLOO on multi-turn TextWorld tasks across model sizes.
Task / Model Base model RLOO PPO
w2-o3-q4 / Qwen-1.5B $0.15$ $0.51_{\uparrow 0.36}$ $0.88_{\uparrow 0.73}$
w4-o6-q8 / Qwen-1.5B $0.01$ $0.0$ $0.59_{\uparrow 0.58}$
w2-o3-q4 / Qwen-7B $0.65$ $0.97_{\uparrow 0.32}$ $0.98_{\uparrow 0.33}$
w4-o6-q8 / Qwen-7B $0.28$ $0.47_{\uparrow 0.21}$ $0.72_{\uparrow 0.44}$
We also examine GRPO, another heuristic-based method, implemented in §5.3 for SWE-Gym tasks. While GRPO does not naturally apply to multi-turn training, SWE-Gym’s heavyweight environment makes GRPO computationally advantageous, nullifying advantages of other methods given the lack of dense reward functions for SWE-Gym. Therefore, we draw parallels between PPO and GRPO as biased algorithms only under sparse reward settings (only final rewards given per trajectory). The improvements from GRPO in Table 6, together with PPO’s results in Table 8, demonstrate the effectiveness of biased methods in multi-turn RL.
@@ 7 Reward
We experiment with different reward densities on TextWorld tasks using PPO and RLOO. Our reward density configurations leverage TextWorld’s built-in reward functions: (1) sparse rewards, provided only upon task completion, and (2) dense rewards, provided at intermediate steps. We quantify reward density as the average number of steps between reward signals and larger values indicate sparser rewards. We evaluate the tw-simple tasks [^4] https://textworld.readthedocs.io/en/stable/textworld.challenges.simple.html from TextWorld, training on 3,000 episodes with different random seeds. Full task specifications and training details appear in Appendices B.3 and C.3.
@@ 7.1 What reward signals are needed for multi-turn RL training?
Dense rewards significantly improve multi-turn RL performance, with optimal density varying by algorithm. Table 9 shows differential effects on PPO and RLOO. PPO benefits most from the densest rewards (Dense 2), achieving 58% success versus 41% with sparse rewards. RLOO shows robust performance across dense schemes, achieving 55% success with both Dense 1 and Dense 2 configurations, suggesting its unbiased gradient estimates are less sensitive to reward density.
The key insight is that reward density should align with the optimization algorithm. While dense rewards enable faster convergence—evidenced by substantial gains across both algorithms—their effectiveness depends on signal quality. Poorly designed intermediate rewards can provide misleading guidance that impairs learning.
@@ 8 The Recipe and Conclusion
Our systematic investigation across environment, policy, and reward pillars yields a practical recipe for multi-turn agentic RL.
Environment: Start training on simpler environments as agents may develop transferable skills that generalize to complex settings. Object complexity proves more challenging than spatial complexity, suggesting that curriculum design should prioritize object manipulation skills. While single-task training provides decent cross-task generalization, mixed-task training yields superior robustness.
Policy: Good imitation learning priors dramatically reduce RL sample complexity while maintaining comparable performance. An optimal balance exists between demonstration and RL data that maximizes both task-specific accuracy and generalization. Biased algorithms (PPO, GRPO) outperform unbiased alternatives (RLOO) in multi-turn settings, with performance gaps widening in complex environments.
Reward: Dense rewards significantly improve multi-turn RL performance, with optimal density varying by algorithm—PPO benefits most from dense feedback while RLOO shows robustness across reward schemes.
Through systematic hyperparameter tuning, we identify a recipe that works across TextWorld, ALFWorld, and SWE-Gym, providing a starting point for practitioners. This work establishes that multi-turn RL requires fundamental rethinking beyond single-turn optimization. By releasing our framework and experimental artifacts, we aim to accelerate research in agentic AI systems capable of extended real-world interaction, providing both the empirical foundation and practical guidelines for developing truly autonomous agents.
@@ C.3 Additional training details for §7
Appendix D Use of LLMs

### v2

@@ 6.1 How does prior model policy influence multi-turn RL training?
We distinguish between establishing a prior (via SFT) and refining it through continual learning (via multi-turn online RL). This two-stage approach mirrors real-world deployment where agents learn from demonstrations before online learning. We find that a strong SFT prior dramatically accelerates RL efficiency. Training on golden solutions from TextWorld tasks w2-o3-q4 reveals that we can match the performance of extensive RL training with a fraction of the data. As shown in Table 9, a model initialized with just 60 demonstrations and refined over 400 RL episodes achieves 85% success, nearly matching the 88% achieved by pure RL training that requires 5000 episodes.
@@
Finally, we explore whether priors transfer across distinct environments by training SFT models on ALFWorld (3553 samples) to initialize TextWorld agents, and vice versa. We find that cross-domain priors actively harm multi-turn RL training. Initializing with a different domain causes rapid policy collapse, likely because the specific behavioral biases and action distributions learned from the source environment conflict with the dynamics of the target environment, effectively destabilizing the policy optimization.
@@ 6.2 How do RL algorithms impact multi-turn RL training?
Understanding whether performance gains stem from our multi-turn formulation or specific algorithmic choices is crucial for establishing the generalizability of our approach. We compare PPO (heuristic policy gradient method with value function bootstrapping) against RLOO (unbiased policy gradient estimator), as well as recent variants like Reinforce++ and GRPO, to isolate the contributions of our token-level credit assignment from algorithmic design decisions (heuristics-harmful).
Both PPO and RLOO achieve improvements over base models, showing that the performance gains stem from multi-turn formulation rather than PPO-specific heuristics. Table 10 shows that PPO achieves 88% success on w2-o3-q4 versus RLOO’s 51%. In contrast, Reinforce++ and GRPO induce negligible gains over the base model. This gap widens on harder tasks w4-o6-q8: PPO maintains 59% success while RLOO, Reinforce++, and GRPO all suffer collapse for the 1.5B model.
Model scaling helps close the gap on simpler tasks, and PPO maintains the best algorithm in complex reasoning. At 7B parameters, both PPO and RLOO converge on simple w2-o3-q4 tasks ($\approx$ 98%), but Reinforce++ and GRPO still lag behind (72% and 79% respectively. On the harder w4-o6-q8 tasks, PPO dominates with 72% success compared to RLOO’s 47% and GRPO’s 36%. These results confirm that the performance gains are not due to PPO heuristics, evidenced by RLOO’s consistent improvements across tasks, but they do highlight that PPO is significantly more sample-efficient and robust than Reinforce++ and GRPO in this domain.
Table 10: Comparison of different RL algorithms, including PPO, RLOO, Reinforce++, and GRPO on multi-turn TextWorld tasks across model sizes.
Task / Model Base model RLOO PPO Reinforce++ GRPO
w2-o3-q4 / Qwen-1.5B $0.15$ $0.51_{\uparrow 0.36}$ $0.88_{\uparrow 0.73}$ $0.18_{\uparrow 0.03}$ $0.18_{\uparrow 0.03}$
w4-o6-q8 / Qwen-1.5B $0.01$ $0.0$ $59_{\uparrow 0.58} 0.0$0.0 $0.02_{\uparrow 0.01}$
w2-o3-q4 / Qwen-7B $0.65$ $0.97_{\uparrow 0.32}$ $0.98_{\uparrow 0.33}$ $0.72_{\uparrow 0.07}$ $0.79_{\uparrow 0.14}$
w4-o6-q8 / Qwen-7B $0.28$ $0.47_{\uparrow 0.21}$ $0.72_{\uparrow 0.44}$ $0.33_{\uparrow 0.05}$ $0.36_{\uparrow 0.08}$
We also examine GRPO in §5.3 for SWE-Gym tasks. While GRPO underperforms on the TextWorld tasks in Table 10, SWE-Gym’s environment makes GRPO computationally advantageous. Therefore, we draw parallels between PPO and GRPO as biased algorithms only under sparse reward settings (only final rewards per trajectory). The improvements from GRPO in Table 8, together with PPO’s results in Table 10, show the effectiveness of biased methods in multi-turn RL.
@@ 7 Reward
We investigate the impact of reward design across two distinct domains: the text-based game TextWorld and the software engineering environment SWE-Gym. For TextWorld, we manipulate reward density using built-in functions, comparing sparse rewards (task completion only) against dense rewards (intermediate steps). We evaluate these on the tw-simple tasks [^2] https://textworld.readthedocs.io/en/stable/textworld.challenges.simple.html using PPO and RLOO over 3,000 episodes. For SWE-Gym, we introduce a more complex evaluation of reward source and granularity. We compare verified rewards derived from ground-truth execution (both binary task completion and granular unit test ratios) against model-based rewards synthesized by judges (CodeRM-8B (coderm) and GPT-4.1). Given the long horizons of coding tasks, we employ GRPO for stability. Full specifications for both environments and reward configurations are detailed in Appendices B.3 and C.3.
@@ 7.1 What reward signals are needed for multi-turn RL training?
Dense rewards significantly improve multi-turn RL performance, with optimal density varying by algorithm. As shown in Table 11, PPO benefits from more frequent feedback, with success rates jumping from 41% under sparse rewards to 58% with the densest signal (Dense 2). In contrast, RLOO maintains steady performance (55%) across both dense configurations, suggesting its unbiased gradient estimates are less sensitive to reward density.
The key insight here is that reward density should be calibrated to the optimization strategy. While dense signals can accelerate convergence, their utility depends on the quality of that signal. If intermediate rewards are poorly designed, they risk introducing noise that misguides the agent.
7.2 Verified vs. Model-Based Rewards in Complex Tasks
While TextWorld provides a controlled setting for analyzing reward density, real-world applications such as software engineering often involve complex and long-horizon reasoning where defining reward density is more challenging. To address this, we extend our analysis to the coding domain using SWE-Gym. We investigate two critical questions: (1) Does the granularity of verified feedback matter? and (2) Can model-based judges substitute for verifiers?
The Necessity of Fine-Grained Verifier Rewards.
We first compare two implementations of execution-based feedback: a sparse binary verified reward (where the agent receives +1 only upon passing the full test suite) and a dense ratio-based verified reward (proportional to the percentage of unit tests passed). As is shown in Table 12, fine-grained verifier rewards are essential for more complex tasks. With sparse binary rewards, the agent achieves only a 4.2% success rate, which is a negligible improvement over the base model’s 4%. However, providing dense, ratio-based feedback boosts performance to 22%. This confirms that in complex domains like software engineering, the “all-or-nothing” signal of a final test suite is too sparse to guide the policy, whereas fine-grained rew for passing individual unit tests provides the necessary gradient for learning.
Verifiers vs. Model-Based Judges.
We also explore whether model-based rewards can serve as an effective alternative to verifiers. We employ two model-based judges capable of code generation: CodeRM-8B (coderm) and GPT-4.1, to generate synthetic test suites and evaluate agent’s code generation progress. While model-based rewards show some improvement over the base model, they significantly underperform verified rewards. As shown in Table 12, rewards shaped by CodeRM-8B and GPT-4.1 achieve success rates of 7.2% and 9.3% respectively. Although these model-based judges provide some useful signal, they lag behind verified ratio rewards (22%) by a wide margin. Consequently, for practitioners, the most robust recipe remains relying on execution-based unit tests rather than model-based proxies.
Table 12: Performance of GRPO on SWE-Gym tasks with different reward signals. We compare the base model against verified rewards (binary vs. ratio) and model-based rewards (CodeRM-8B vs. GPT-4.1).
Metric Base Model Verified (Binary) Verified (Ratio)
Success Rate $0.04$ $0.042_{\uparrow 0.002}$ $0.22_{\uparrow 0.18}$
Model-Based Rewards Judge (CodeRM-8B) Judge (GPT-4.1)
Success Rate $0.04$ $072_{\uparrow 0.032} 0.0$$093_{\uparrow 0.053}$
@@ 8 The Recipe and Conclusion
Our systematic investigation across environment, policy, and reward pillars yields a practical recipe for multi-turn agentic RL, validated across TextWorld, ALFWorld, and SWE-Gym.
Environment: Curriculum matters. Start training on simpler environments or tasks with lower verification complexity. We find that agents develop reusable skills (e.g., state tracking) in easier settings that generalize effectively to complex, object-heavy, or long-horizon tasks.
Policy: Prioritize stability and priors. Biased, stabilized algorithms (PPO, GRPO) significantly outperform unbiased estimators (RLOO) and naive gradients (Reinforce++) in multi-turn settings. Furthermore, initializing with a strong SFT prior dramatically reduces RL sample complexity, offering a more efficient path to convergence than pure RL.
Reward: Granularity beats approximation. Dense rewards are essential for complex reasoning. However, our results show that verified execution feedback (e.g., unit test pass rates) is far superior to model-based surrogates. Practitioners should prioritize engineering granular, execution-based signals over training reward models.
By systematically identifying this recipe: Curriculum + Stabilized Policy + Verified Dense Reward, we provide the empirical foundation and practical guidelines for developing truly autonomous agents capable of extended real-world interaction.
@@ C.3 Additional training details for §7
Appendix D Training curves for PPO versus RLOO policies on TextWorld tasks
Figure 6: The training reward curve for Qwen-2.5-1.5B model trained using PPO and RLOO on TextWorld w2-o3-q4 tasks.
Figure 7: The training reward curve for Qwen-2.5-7B model trained using PPO and RLOO on TextWorld w4-o6-q8 tasks.
Appendix E Use of LLMs